OpenStudy (anonymous):

translate and rotate the hyperbola 2 * x ^ 2 +6 * x * y +2 * y = 10 * (x + y +1)

OpenStudy (anonymous):

This is the Equation \[x ^{2} + 3*x*y -5*x -4*y -5 =0\]

OpenStudy (anonymous):

hyperbola equation== \[x^{2} \div a ^{2} - y^{2} \div b ^{2} =1\]

OpenStudy (anonymous):

will be how to think

OpenStudy (experimentx):

for translation x = X - h y = Y - k

OpenStudy (experimentx):

For rotation \[ x = X \cos \theta - Y \sin \theta \\ y = Y \cos \theta + X \sin \theta \]

OpenStudy (anonymous):

ok, but what i do first? translate or rotate?

OpenStudy (experimentx):

do you want to reduce it to standard form?? \[ \frac {x^2}a - \frac {y^2} b = 1 \] ??

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

that's it, but to get this form i need to translate and rotate

OpenStudy (experimentx):

It's indeed a hyperbola i guess translation first http://www.wolframalpha.com/input/?i=plot+2+*+x+^+2+%2B6+*+x+*+y+%2B2+*+y+%3D+10+*+%28x+%2B+y+%2B1%29 then rotation

OpenStudy (experimentx):

Let me check first ... i'm not sure i haven't done this for very long time

OpenStudy (anonymous):

ok xD

OpenStudy (experimentx):

I guess rotation first ... to eliminate the homogeneous form xy http://staff.argyll.epsb.ca/jreed/math30p/conics/standard.htm

OpenStudy (experimentx):

this is going to be long work !!! do you have some kind of software for this type of job??

OpenStudy (anonymous):

i use woframalpha sometimes, but i'll need to do this by hand heheh

OpenStudy (anonymous):

ok, i'll start rotating first

OpenStudy (anonymous):

thanks a lot!

OpenStudy (experimentx):

best of luck ... when many terms are involved, my sight get worse ...